# A mixed logic with binary operators

### v1

@@ 4. The class $\operatorname{\mathbf{V}}(\mathbf{wMIA})$
Let $\operatorname{{\mathsf{Cf}}}(\mathfrak{A})=\langle\operatorname{Ult}(A),Q_{f},S_{g}\rangle$ be the canonical frame of $\mathfrak{A}$. Since $\operatorname{{\mathsf{Cf}}}(\mathfrak{A})$ is a wMIA frame, we construct its special frame following the method from subsection 3.3 (see Figure 7). To this end, let us consider an isomorphic copy $\operatorname{{\mathsf{Cf}}}^{\prime}(\mathfrak{A})=\langle\operatorname{Ult}(A)^{\prime},Q^{\prime}_{f},S^{\prime}_{g}\rangle$ of $\operatorname{{\mathsf{Cf}}}(\mathfrak{A})$ where $\operatorname{Ult}(A)$ and $\operatorname{Ult}^{\prime}(A)$ are disjoint sets. We put $\underline{\operatorname{Ult}A}\coloneqq\operatorname{Ult}(A)\cup\operatorname{Ult}(A)^{\prime}$, and define $\underline{R}\subseteq\underline{\operatorname{Ult}A}^{3}$ and its complement as before:
@@
$s(a)\coloneqq\{\mathscr{U}\in\operatorname{Ult}(A):a\in\mathscr{U}\}\cup\{\mathscr{U}^{\prime}\in\operatorname{Ult}(A)^{\prime}:a\in\mathscr{U}\}\,.$
@@
Let $\mathscr{U}^{\epsilon}\in s(f(a,b))$. By definition, we have that $f(a,b)\in\mathscr{U}$. Thus, there exist $\mathscr{U}_{a},\mathscr{U}_{b}\in\mathscr{U}(A)$ such that $a\in\mathscr{U}_{a}$, $b\in\mathscr{U}_{b}$ and $f[\mathscr{U}_{a}\times\mathscr{U}_{b}]\subseteq\mathscr{U}$, [^8] See, e.g., (Jónsson and Tarski, 1951). i.e., $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in Q_{f}$. By construction, $\langle\mathscr{U}^{\prime}_{a},\mathscr{U}^{\epsilon},\mathscr{U}_{b}\rangle\in\underline{R}$. It follows that $(s(a)\times\{\mathscr{U}^{\epsilon}\}\times s(b))\cap\underline{R}\neq\emptyset$ and therefore $\mathscr{U}^{\epsilon}\in\langle{\underline{R}\rangle}(s(a),s(b))$.
For the other direction, let $\mathscr{U}^{\epsilon}\in\langle{\underline{R}\rangle}(s(a),s(b))$, i.e., $(s(a)\times\{\mathscr{U}^{\epsilon}\}\times s(b))\cap\underline{R}\neq\emptyset$. Then, there exist $\mathscr{U}_{a},\mathscr{U}_{b}\in\operatorname{Ult}(A)$ such that $\mathscr{U}^{\epsilon}_{a}\in s(a)$, $\mathscr{U}^{\epsilon}_{b}\in s(b)$ and $\langle\mathscr{U}^{\epsilon}_{a},\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon}_{b}\rangle\in\underline{R}$. By construction, $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in S_{g}$ or $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in Q_{f}\cap-S_{g}$. In both cases, $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in Q_{f}$, which implies that $f[\mathscr{U}_{a}\times\mathscr{U}_{b}]\subseteq\mathscr{U}$ and therefore $f(a,b)\in\mathscr{U}$. By definition, we obtain $\mathscr{U}^{\epsilon}\in s(f(a,b))$.
@@
For the reverse inclusion, assume that $\mathscr{U}^{\epsilon}\in[[\underline{R}]](s(a),s(b))$, that is $s(a)\times\{\mathscr{U}^{\epsilon}\}\times s(b)\subseteq\underline{R}$. Let $\mathscr{U}_{a},\mathscr{U}_{b}\in\operatorname{Ult}(A)$ be such that $a\in\mathscr{U}_{a}$ and $b\in\mathscr{U}_{b}$. Then, $\langle\mathscr{U}_{a},\mathscr{U}^{\epsilon},\mathscr{U}_{b}\rangle\in\underline{R}$. Suppose that $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\notin S_{g}$. By construction we have that $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in Q_{f}$ but then $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\notin\underline{R}$ and $\langle\mathscr{U}^{\prime}_{a},\mathscr{U}^{\prime},\mathscr{U}^{\prime}_{b}\rangle\notin\underline{R}$. This contradicts the fact that $s(a)\times\{\mathscr{U}^{\epsilon}\}\times s(b)\subseteq\underline{R}$, and thus, $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in S_{g}$. Since $\mathscr{U}_{a},\mathscr{U}_{b}$ are arbitrary ultrafilters that contain $a$ and $b$, respectively, we obtain $g(a,b)\in\mathscr{U}$. We conclude that $\mathscr{U}^{\epsilon}\in s(g(a,b))$.
@@
The following corollary is a counterpart of (Düntsch et al., 2017, Theorem 8.5).
@@ 6. Applications to b-algebras
We will show now how the results obtained in this paper can be applied to betweenness algebras introduced and studied in Düntsch et al. (2023). We begin by recalling the key concepts from the paper.
@@
In (Düntsch et al., 2023, Example 8.4) we demonstrated that there is a b-algebra that cannot be embedded into the complex algebra of any b-frame. However, as a consequence of Theorem 5.1 we have the following
@@
As mentioned before, this answers problem (1) from (Düntsch et al., 2023, Section 9). We know from Example 8.4 of the paper that the ternary relation $\underline{R}$ from Theorem 5.1 cannot be a betweenness relation, yet we may ask about its general properties. On the positive side, we have that $\underline{R}$ satisfies (BT1)
@@
In the first case, by (Düntsch et al., 2023, Lemma 7.2) we obtain that the triple $\langle\mathscr{U}_{3},\mathscr{U}_{2},\mathscr{U}_{1}\rangle$ is in $S_{g}$, so
@@
By the construction of $\underline{R}$, the only triples of the form $\langle\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon}\rangle$ that are in $\underline{R}$ can only be in the cells of triples in $S_{g}$. However, by (Düntsch et al., 2023, Theorem 6.10), if $S_{g}(\mathscr{U},\mathscr{U},\mathscr{U})$, then $\mathscr{U}$ must be a principal ultrafilter. Therefore, if $\underline{R}(\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon})$, then $\mathscr{U}^{\epsilon}$ is either a principal ultrafilter or a copy of a principal ultrafilter. ∎
Proposition 6.6.
@@
The proof is based on (Düntsch et al., 2023, Example 8.6). We take the 3-frame $\mathfrak{N}=\langle\omega,B_{\leqslant}\rangle$ where $B_{\leqslant}$ is the betweenness relation obtained from the standard binary order $\leqslant$ on the set $\omega$ of natural numbers. Let $2^{\omega}$ be the set of all sets of natural numbers. We consider the complex algebra $\operatorname{{\mathsf{Cm}}}(\mathfrak{N})=\langle 2^{\omega},Q_{\langle{B_{\leqslant}\rangle}},S_{[[B_{\leqslant}]]}\rangle$, its canonical frame $\operatorname{{\mathsf{Cf}}}(\operatorname{{\mathsf{Cm}}}(\mathfrak{N}))=\langle\operatorname{Ult}(2^{\omega}),Q_{\langle{B_{\leqslant}\rangle}},S_{[[B_{\leqslant}]]}\rangle$, and its special frame $\langle\underline{\operatorname{Ult}2^{\omega}},\underline{R}\rangle$.
For any triple of free ultrafilters $\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}$ it is the case that $Q_{\langle{B_{\leqslant}\rangle}}(\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3})$. Recall that also $-S_{[[B_{\leqslant}]]}(\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3})$ for any such triple. Therefore $\underline{R}(\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}^{\prime})$, since
@@
But $-\underline{R}(\mathscr{U}_{1},\mathscr{U}_{1},\mathscr{U}_{2})$, as $Q_{\langle{B_{\leqslant}\rangle}}(\mathscr{U}_{1},\mathscr{U}_{1},\mathscr{U}_{2})$ and $-S_{[[B_{\leqslant}]]}(\mathscr{U}_{1},\mathscr{U}_{1},\mathscr{U}_{2})$. Consequently, (BT2) fails for the special frame $\langle\underline{\operatorname{Ult}2^{\omega}},\underline{R}\rangle$.
@@
In Düntsch et al. (2023) we also considered two more axioms for the betweenness:
@@
The reader will easily check that the special model $\langle\underline{\operatorname{Ult}2^{\omega}},\underline{R}\rangle$ does not satisfy any of the two axioms.
Acknowledgments

### v2

@@ 4. The class $\operatorname{\mathbf{V}}(\mathbf{wMIA})$
Let $\operatorname{{\mathsf{Cf}}}(\mathfrak{A})=\langle\operatorname{Ult}A,Q_{f},S_{g}\rangle$ be the canonical frame of $\mathfrak{A}$. Since $\operatorname{{\mathsf{Cf}}}(\mathfrak{A})$ is a wMIA frame, we construct its special frame following the method from subsection 3.3 (see Figure 7). To this end, let us consider an isomorphic copy $\operatorname{{\mathsf{Cf}}}^{\prime}(\mathfrak{A})=\langle\operatorname{Ult}A^{\prime},Q^{\prime}_{f},S^{\prime}_{g}\rangle$ of $\operatorname{{\mathsf{Cf}}}(\mathfrak{A})$ where $\operatorname{Ult}A$ and $\operatorname{Ult}A^{\prime}$ are disjoint sets. We put $\underline{\operatorname{Ult}A}\coloneqq\operatorname{Ult}A\cup\operatorname{Ult}A^{\prime}$, and define $\underline{R}\subseteq\underline{\operatorname{Ult}A}^{3}$ and its complement as before:
@@
$s(a)\coloneqq\{\mathscr{U}\in\operatorname{Ult}A:a\in\mathscr{U}\}\cup\{\mathscr{U}^{\prime}\in\operatorname{Ult}A^{\prime}:a\in\mathscr{U}\}.$
@@
Let $\mathscr{U}^{\epsilon}$ denote either the ultrafilter $\mathscr{U}$ or its copy $\mathscr{U}^{\prime}$. Assume that $\mathscr{U}^{\epsilon}\in s(f(a,b))$. By the definition, we have that $f(a,b)\in\mathscr{U}$. Thus, there exist $\mathscr{U}_{a},\mathscr{U}_{b}\in\operatorname{Ult}A$ such that $a\in\mathscr{U}_{a}$, $b\in\mathscr{U}_{b}$ and $f[\mathscr{U}_{a}\times\mathscr{U}_{b}]\subseteq\mathscr{U}$, [^8] See, e.g., [11]. i.e., $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in Q_{f}$. By construction, $\langle\mathscr{U}^{\prime}_{a},\mathscr{U}^{\epsilon},\mathscr{U}_{b}\rangle\in\underline{R}$. It follows that $(s(a)\times\{\mathscr{U}^{\epsilon}\}\times s(b))\cap\underline{R}\neq\emptyset$ and therefore $\mathscr{U}^{\epsilon}\in\langle{\underline{R}\rangle}(s(a),s(b))$.
For the other direction, let $\mathscr{U}^{\epsilon}\in\langle{\underline{R}\rangle}(s(a),s(b))$, i.e., $(s(a)\times\{\mathscr{U}^{\epsilon}\}\times s(b))\cap\underline{R}\neq\emptyset$. Then, there exist $\mathscr{U}_{a},\mathscr{U}_{b}\in\operatorname{Ult}A$ such that $\mathscr{U}^{\epsilon}_{a}\in s(a)$, $\mathscr{U}^{\epsilon}_{b}\in s(b)$ and $\langle\mathscr{U}^{\epsilon}_{a},\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon}_{b}\rangle\in\underline{R}$. By construction, $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in S_{g}$ or $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in Q_{f}\cap-S_{g}$. In both cases, $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in Q_{f}$, which implies that $f[\mathscr{U}_{a}\times\mathscr{U}_{b}]\subseteq\mathscr{U}$ and therefore $f(a,b)\in\mathscr{U}$. By definition, we obtain $\mathscr{U}^{\epsilon}\in s(f(a,b))$.
@@
For the reverse inclusion, assume that $\mathscr{U}^{\epsilon}\in[[\underline{R}]](s(a),s(b))$, that is $s(a)\times\{\mathscr{U}^{\epsilon}\}\times s(b)\subseteq\underline{R}$. Let $\mathscr{U}_{a},\mathscr{U}_{b}\in\operatorname{Ult}A$ be such that $a\in\mathscr{U}_{a}$ and $b\in\mathscr{U}_{b}$. Then, $\langle\mathscr{U}_{a},\mathscr{U}^{\epsilon},\mathscr{U}_{b}\rangle\in\underline{R}$. Suppose that $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\notin S_{g}$. By construction we have that $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in Q_{f}$ but then $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\notin\underline{R}$ and $\langle\mathscr{U}^{\prime}_{a},\mathscr{U}^{\prime},\mathscr{U}^{\prime}_{b}\rangle\notin\underline{R}$. This contradicts the fact that $s(a)\times\{\mathscr{U}^{\epsilon}\}\times s(b)\subseteq\underline{R}$, and thus, $\langle\mathscr{U}_{a},\mathscr{U},\mathscr{U}_{b}\rangle\in S_{g}$. Since $\mathscr{U}_{a},\mathscr{U}_{b}$ are arbitrary ultrafilters that contain $a$ and $b$, respectively, we obtain $g(a,b)\in\mathscr{U}$. We conclude that $\mathscr{U}^{\epsilon}\in s(g(a,b))$.
@@
The following corollary is a counterpart of [5, Theorem 8.5].
@@ 6. Applications to b-algebras
We will show now how the results obtained in this paper can be applied to betweenness algebras introduced and studied in [6]. We begin by recalling the key concepts from the paper.
@@
In [6, Example 8.4] we demonstrated that there is a b-algebra that cannot be embedded into the complex algebra of any b-frame. However, as a consequence of Theorem 5.1 we have the following
@@
As mentioned before, this answers problem (1) from [6, Section 9]. We know from Example 8.4 of the paper that the ternary relation $\underline{R}$ from Theorem 5.1 cannot be a betweenness relation, yet we may ask about its general properties. On the positive side, we have that $\underline{R}$ satisfies (BT1)
@@
In the first case, by [6, Lemma 7.2] we obtain that the triple $\langle\mathscr{U}_{3},\mathscr{U}_{2},\mathscr{U}_{1}\rangle$ is in $S_{g}$, so
@@
By the construction of $\underline{R}$, the only triples of the form $\langle\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon}\rangle$ that are in $\underline{R}$ can only be in the cells of triples in $S_{g}$. However, by [6, Theorem 6.10], if $S_{g}(\mathscr{U},\mathscr{U},\mathscr{U})$, then $\mathscr{U}$ must be a principal ultrafilter. Therefore, if $\underline{R}(\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon},\mathscr{U}^{\epsilon})$, then $\mathscr{U}^{\epsilon}$ is either a principal ultrafilter or a copy of a principal ultrafilter. ∎
Proposition $6.$
@@
The proof is based on [6, Example 8.6]. We take the 3-frame $\mathfrak{N}=\langle\omega,B_{\leqslant}\rangle$ where $B_{\leqslant}$ is the betweenness relation obtained from the standard binary order $\leqslant$ on the set $\omega$ of natural numbers. Let $2^{\omega}$ be the set of all sets of natural numbers. We consider the complex algebra $\operatorname{{\mathsf{Cm}}}(\mathfrak{N})=\langle 2^{\omega},Q_{\langle{B_{\leqslant}\rangle}},S_{[[B_{\leqslant}]]}\rangle$, its canonical frame $\operatorname{{\mathsf{Cf}}}(\operatorname{{\mathsf{Cm}}}(\mathfrak{N}))=\langle\operatorname{Ult}2^{\omega},Q_{\langle{B_{\leqslant}\rangle}},S_{[[B_{\leqslant}]]}\rangle$, and its special frame $\langle\underline{\operatorname{Ult}2^{\omega}},\underline{R}\rangle$.
For any triple of free ultrafilters $\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}$ it is the case that $Q_{\langle{B_{\leqslant}\rangle}}(\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3})$. Recall that also $S_{[[B_{\leqslant}]]}(\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3})$ for any such triple. Therefore $\underline{R}(\mathscr{U}_{1},\mathscr{U}_{2},\mathscr{U}_{3}^{\prime})$, since
@@
But $\underline{R}(\mathscr{U}_{1},\mathscr{U}_{1},\mathscr{U}_{2})$, as $Q_{\langle{B_{\leqslant}\rangle}}(\mathscr{U}_{1},\mathscr{U}_{1},\mathscr{U}_{2})$ and $S_{[[B_{\leqslant}]]}(\mathscr{U}_{1},\mathscr{U}_{1},\mathscr{U}_{2})$. Consequently, (BT2) fails for the special frame $\langle\underline{\operatorname{Ult}2^{\omega}},\underline{R}\rangle$.
@@
In [6] we also considered two more axioms for the betweenness:
@@
The reader will easily check that the special model $\langle\underline{\operatorname{Ult}2^{\omega}},\underline{R}\rangle$ does not satisfy any of them.
Funding